The sum of 4 times mikes age and 3 times susan's age is 79. mike is 1 year more than 3 times as old as susan. find each of their ages.
Answer by Maths68(1474)

Let
Present age of mike = m
Present age of Susan = s
Given
======
The sum of 4 times mikes age and 3 times susan's age is 79
4m+3s=79..............(1)
Mike is 1 year more than 3 times as old as Susan
m=3s+1................(2)
put the value of m from (2) to (1)
m=3s+1
4(3s+1)+3s=79
12s+4+3s=79
15s+4=79
15s=79-4
15s=75
15s/15=75/15
s=5
put the value of s in (2)
m=3s+1................(2)
m=3(5)+1
m=15+1
m=16

Present age of mike = m = 16 years old
Present age of Susan = s = 5 years old